OpenStudy (givingtree):

Please help me on how to solve for the Future value. Morgan wants to purchase a home in six years. He will contribute $3500 each year to a savings account with 3.26% interest, compounded quarterly. What is the future value of this investment, when Maurice needs to make a down payment?

OpenStudy (givingtree):

@mathmale @Awolflover1 @Luigi0210 @retirEEd @Jaynator495 Future value help?

OpenStudy (retireed):

Does he contribute the $3500 all at once or smaller amounts during each year?

OpenStudy (givingtree):

I think he will contribute $3500 each year. Would the interest apply though?

OpenStudy (givingtree):

\[FV = PV(1+i) ^{nt}\]

OpenStudy (givingtree):

Pretty sure that is the formula.

OpenStudy (givingtree):

?

OpenStudy (solomonzelman):

\(\large\color{black}{{\bbox[5pt, lightyellow ,border:2px solid black ]{ {\rm F}={\rm I}\times \left(1+\frac{r}{n}\right)^{n\times t} }}}\) where, \(\color{black}{\displaystyle r }\) - percent rate as a decimal, \(\color{black}{\displaystyle n}\) - number of times the interest is compound per year. \(\color{black}{\displaystyle t}\) - number of years. \(\color{black}{\displaystyle {\rm I}}\) - initial investment. \(\color{black}{\displaystyle {\rm F}}\) - the resulting value after the investment.

OpenStudy (givingtree):

Thank you! I'll plug in the numbers and solve, can i check back with you to see if i am correct?

OpenStudy (solomonzelman):

Sure:) (If I'm online)

OpenStudy (givingtree):

Thanks!